OpenStudy (anonymous):

(6-2x)^2-144=0 Can you help me understand how you solve this and why you solve it in such a way.

OpenStudy (compassionate):

This is a quadratic equation. You can solve it using the quadratic formula. Check this out. http://www.purplemath.com/modules/quadform.htm

OpenStudy (icecreamyummy):

X=-9,-3

OpenStudy (campbell_st):

a simple step is to add 144 to both sides of the equation \[(6 - 2x)^2 = 144\] then take the square root of both sides \[6 - 2x = \pm 12\] which can become 6 - 2x = 12 and 6 - 2x = -12 so solve both equations to get the solutions to the equations.

OpenStudy (anonymous):

http://i.gyazo.com/7bf6fc045a918679b3b59fc6c1f78c77.png how do you know if the 9 and the 3 is possitive or negative